# Biodenoising: Animal Vocalization Denoising without Access to Clean Data

### v2

@@ II Relation with previous work
The lack of clean training data presents a challenge for model development, similar to unsupervised speech enhancement [14, 15, 16, 17, 18, saijo2023self]. One way to approach this task is training on mixtures of noisy targets and noise, known as ‘noisy target training’ [17]. Here we prove that leveraging existing speech enhancement models to obtain pseudo-clean targets is better than using noisy targets.
To our best knowledge, besides generic denoising methods such as ‘noisereduce’ [9], there are no state of the art methods that we can use across multiple species and environments [19, 20, 21]. Instead, we train a model that is tested for generalization capabilities on a non-overlapping benchmark. We take inspiration in recent work showing that self-supervised speech models offer good priors for various bioacoustics classification tasks [22, 23, 24] and that classifier models trained on bird signals perform well when used for other taxa [25]. These recent findings suggest that cross-taxa similarities in signal characteristics and recording environments may allow for generalization.
@@ IV Methods
We explore whether the parameters of the speech denoising method are good priors when learning . This formulation is similar to the teacher-student domain adaptation approach [16, saijo2023self] for a single update. Our research hypothesis is that the estimations obtained from training with pseudo-clean targets have better quality than the pseudo-clean targets themselves and than the estimations from the ‘noisy target’ baseline [19], for unseen contexts.

### v3

@@ II Relation with previous work
The lack of clean training data presents a challenge for model development, similar to unsupervised speech enhancement [14, 15, 16, 17, 18, 19]. One way to approach this task is training on mixtures of noisy targets and noise, known as ‘noisy target training’ [17]. Here we prove that leveraging existing speech enhancement models to obtain pseudo-clean targets is better than using noisy targets.
To our best knowledge, besides generic denoising methods such as ‘noisereduce’ [9], most of the methods are not evaluated across multiple species and environments [20, 21, 22, 23]. Instead, we train a model that is tested for generalization capabilities on a non-overlapping benchmark. We take inspiration in recent work showing that self-supervised speech models offer good priors for various bioacoustics classification tasks [24, 25, 26] and that classifier models trained on bird signals perform well when used for other taxa [27]. These recent findings suggest that cross-taxa similarities in signal characteristics and recording environments may allow for generalization.
@@ IV Methods
We explore whether the parameters of the speech denoising method are good priors when learning . This formulation is similar to the teacher-student domain adaptation approach [16, 19] for a single update. Our research hypothesis is that the estimations obtained from training with pseudo-clean targets have better quality than the pseudo-clean targets themselves and than the estimations from the ‘noisy target’ baseline [20], for unseen contexts.